Cal Bank believes the US dollar will appreciate over the next five days from GHS 4.48 to GHS 4.50. The following annual interest rates apply:
Currency. Lending Rate Borrowing Rate
Dollars 7.10% 7.50%
Ghana Cedis 6.80% 7.25%

Cal Bank has the capacity to borrow either GHS 10 million or $5 million. If Cal Bank's forecast is correct, what will its dollar profit be from speculation over the five-days period (assuming it does not use any of its existing consumer deposits to capitalize on its expectations)?
1
Expert's answer
2021-08-10T09:08:06-0400

The speculative profit can be made in the following way :

i) Borrow GHS 10million for 5 days at 7.25% , Amount to be repaid after 5 days = GHS 10 million "\times" (1+0.0725"\times" "\frac{5}{365}" ) = GHS 10,009,932

ii) Convert the GHS 10 million to Dollars to get "\frac{10million}{4.48}" = "\$" 2,232,142.86

iii) Invest this amount for 5 days at 7.10% to get $ 2232142.86 "\times" (1+0.071"\times" "\frac{5}{365}" ) = "\$" 2,234,313.85 after 5 days

iv) After 5 days, get $2,234,313.85 and sell it at 4.50 (if speculation is correct) to get GHS 2234313.85"\times" 4.50 = GHS 10,054,412

v) Repay the GHS borrowing amount of GHS 10,009,932

and make a speculative profit of GHS 10,054,412 - GHS 10,009,932 = GHS 44481

Profit in Dollars = GHS 44481"\/" 4.50 = $9884.62